# A complete invariant for doodles on a 2-sphere

### 17 Jan 2024 v1

@@
Abstract.
We define a complete invariant for doodles on a 2-sphere which takes values in series of chord diagrams of certain type. The coefficients at the diagrams with $n$ chords are finite type invariants of doodles of order at most $2n$.
1. Introduction
The theory of finite type knot invariants may be thought of as a nilpotency theory for knots. Nilpotent and residually nilpotent groups can be studied by embedding them into algebras of power series; most notably, the free group on $n$ letters $x_{1},\ldots,x_{n}$ maps into the algebra of the noncommutative power series in $X_{1},\ldots,X_{n}$ by
@@ 1. Introduction
The Kontsevich integral contains the complete information about the finite type invariants of a knot. One may be tempted to construct universal finite type knot invariants by using the same elementary method that works for the free groups. Namely, one can encode knots by diagrams of some kind and then assign to a knot a formal sum of the subdiagrams of its diagram. In particular, if knots are represented by so-called clasp diagrams [12], the formal sum of the subdiagrams taken modulo diagrams with $n+1$ clasps is, in fact, a universal invariant of order $n$. This construction, however, is no replacement for the Kontsevich integral since the values of this invariant lie in a certain algebra which is rather more involved than the (already uncomputable) algebra of chord diagrams. Essentially, it is a restatement of the original approach of Vassiliev in different terms.
@@
In the present note we consider the set of equivalence classes of curves in $S^{2}$ without triple self-intersections, called “doodles” by Khovanov [^1] The term “doodles” was first used [5] for a different class of curves. Khovanov’s terminology has now superseded that of Fenn and Taylor. in [9]. We show that the sums of subdiagrams give a finite type invariant of curves of doodles in $S^{2}$, which is complete in the sense that it distinguishes non-equivalent curves.
@@
2. Doodles and their arrow diagrams
@@
A doodle in $S^{2}$ is determined up to isotopy by its arrow diagram [^3] It would be more natural to use the term “Gauss diagram”. However, Arnold in [1] uses the term “Gauss diagram” for the underlying chord diagram of the arrow diagram, so we follow the terminology of Merkov. The term “arrow diagram” has also been used in a somewhat different context in [15]., which is a chord diagram with directed chords. The skeleton (that is, the outer circle) of the chord diagram corresponds to the parameter along the curve, the chords connect the inverse images of the crossings. At each crossing the branches of the doodle can be ordered in such a way that the tangent vectors to the first and the second branch form a positive basis of the tangent space to $S^{2}$. Each chord is of the arrow diagram of the doodle is directed from the first branch to the second branch.
Figure 1. A doodle and its arrow diagram
@@
Note that applying a move $R^{\prime}_{1}$ or $R^{\prime}_{2}$ in a way that decreases the number of crossings to a realizable arrow diagram we again obtain a realizable diagram. It follows that the minimal diagram equivalent to a realizable arrow diagram is also realizable. We arrive to the following
Proposition 1.
@@
A function on the set of isotopy classes of doodles that is preserved by the moves $R_{1},R_{2}$ is called an invariant of doodles in $S^{2}$. Proposition 1 implies that any function on the arrow diagrams that is preserved by the moves $R^{\prime}_{1},R^{\prime}_{2}$ gives rise to an invariant of doodles and that each doodle invariant arises in this way.
Write $\mathbf{D}$ for the set of all arrow diagrams and $\mathbf{Q}\mathbf{D}$ for the vector space over $\mathbf{Q}$ generated by this set. The moves $R^{\prime}_{1},R^{\prime}_{2}$ define a set of relations in $\mathbf{Q}\mathbf{D}$ which equate the left-hand side of a move to the right hand side:
@@
the inverse limit of the $I_{i}$. Theorem 1 implies
Corollary 1.
@@
Let $V_{n}$ be the kernel of the projection $\mathcal{A}_{n+1}\to\mathcal{A}_{n}$. The space $\overline{\mathcal{A}}$ is isomorphic to the product of all the $V_{n}$. Another consequence of Theorem 1 is
Corollary 2.
@@
If a basis is chosen in each $V_{n}$, the map $\overline{I}$ becomes a series of chord diagrams whose coefficients are finite type invariants, and the coefficient at any basis diagram in $V_{n}$ is a inite type invariant of order at most $2n$.
In fact, we will see that Theorem 1 holds even if $\mathcal{A}_{n}$ is replaced by a certain quotient of $\mathcal{A}_{n}$ which admits an explicit description. We consider it in the next section. In Section 4 we will show that $I_{n}$ is an invariant of order $2n$.
3. Quiver diagrams and the diagram invariants
@@
Figure 2. Reduction of a chord diagram
Lemma 1.
@@
Proof.
If a chord diagram $C$ has two pairs of adjacent chords, these pairs are either disjoint or they have one chord in common. In both cases, the result of reducing $C$ with respect to both pairs does not depend on the order. It is obvious in the case when the pairs are disjoint; if they are not, inspection shows that the order of reductions does not matter in this case either. ∎
Quiver diagrams [^4] The term “quiver” here has nothing to do with directed graphs and refers to a pack of arrows. are a generalization of arrow diagrams. A quiver diagram is a chord diagram in which the endpoints of the chords are labelled by nonnegative integers in such a way that the sum of the labels for every chord is at least 1. Quiver diagrams in which every chord has one end labelled by 1 and the other end by 0 may be identified with arrow diagrams if we take the direction of each chord to be from 0 towards 1. Define the degree of a quiver diagram to be the sum of its labels. The underlying chord diagram of a quiver diagram is obtained by forgetting the labels of the endpoints.
One may define the reduction of a quiver diagram with two adjacent chords as the reduction of the underlying chord diagram with the following rule for the labels: the labels of the adjacent endpoints of the adjacent chords are added together while the labels of the other chords remain the same, see Figure 3. We say that a quiver diagram is reduced if its underlying chord diagram is reduced. As in Lemma 1, any sequence of reductions of a given quiver diagram always terminates in the same reduced diagram.
Figure 3. Reduction of a quiver diagram
Let $\mathcal{Q}$ be the vector space over $\mathbf{Q}$ spanned by all the quiver diagrams modulo the subspace spanned by following relations:
@@
Define $\mathcal{Q}_{n}$ to be the quotient of $\mathcal{Q}$ by the subspace spanned by the diagrams of degree greater than $n$. For a reduced chord diagram $C$ without isolated chords and at most $n$ chords, let $\mathcal{Q}_{n}^{C}\subset\mathcal{Q}_{n}$ be the subspace spanned by the quiver diagrams whose underlying chord diagram is $C$.
Proposition 2.
@@
Proposition 3.
Given a reduced chord diagram $C$ without isolated chords, mark one endpoint for each chord. The set of all quiver diagrams of degree at most $n$ with underlying chord diagram $C$ and the labels of the marked endpoints all zero form a basis of $\mathcal{Q}_{n}^{C}$.
@@
from where we get
(3) $D(p,q)=-D(p+1,q-1)+D(p+2,q-1)-D(p+3,q-1)+\ldots.$
@@
Rewriting in this way the relations in $\mathcal{Q}_{n}^{C}$ we obtain a tautology so all the diagrams whose marked endpoints are labelled with 0 are linearly independent. ∎
@@
The second part of Theorem 1 is a consequence of the following statement:
Theorem 2.
@@
Figure 4. A pair of intersecting adjacent chords pointing in the same direction and what they might look like in a doodle. The points $B$ and $C$ must be joined in $S^{2}$ by a segment of the doodle containing no other intersection points, which is impossible.
Proof.
Assume that the arrow diagram $X_{D}$ of a minimal doodle $D$ contains a pair of adjacent chords. These chords must point in the same direction since otherwise they could be eliminated by a move $R_{2}$; this would mean that the doodle is not minimal. These chords cannot intersect each other, see Figure 4; the only possible situation when a pair of adjacent chords appears in a minimal doodle diagram is when these chords are parallel as shown in Figure 5 (a); in fact, in this case there may be a cluster with more than two parallel chords with adjacent endpoints as in Figure 5 (b). Such a cluster of parallel chords with adjacent endpoints divides the arrow diagram $X_{D}$ into two parts and no other chord intersects any of the chords of the cluster. The reduction of $X_{D}$ consists in replacing each maximal cluster of $m$ parallel arrows in $X_{D}$ with one chord whose initial point is marked by 00 and the final point by $m$. Denote the resulting reduced quiver diagram by $r(X_{D})$, see Figure 6. If $X_{D}$ is reduced to begin with, we set $r(X_{D})=X_{D}$.
Figure 5. Doodles whose arrow diagrams have adjacent chords
Figure 6. The diagrams $X_{D}$ and $r(X_{D})$.
Assume $D$ is minimal and has at most $n$ crossings. Since each chord in $r(X_{D})$ has one endpoint marked 0, we can choose a basis in $\mathcal{Q}_{n}$ as in Proposition 3 such that $r(X_{D})$ is a basis element. Then, the coefficient of $r(X_{D})$ in $q_{n}(I_{n}(D))$ is equal to 1 since there is precisely one subdiagram that gives $r(X_{D})$; namely, $X_{D}$ itself. It follows that $q_{n}(I_{n}(D))\neq 0$.
Now, let $D_{1}\neq D_{2}$ be two minimal doodles with at most $n$ crossings. Without loss of generality, assume that $D_{1}$ has at least as many crossings as $D_{2}$. Then, if the the underlying chord diagrams of the arrow diagrams $X_{D_{1}}$ and $X_{D_{2}}$ are different, $r(X_{D_{1}})\neq r(X_{D_{2}})$ and we can choose a basis for $\mathcal{Q}_{n}$ in such a way that both $r(X_{D_{1}})$ and $r(X_{D_{2}})$ are basis elements. Then, the coefficient of $r(X_{D_{1}})$ in $q_{n}(I_{n}(D_{1}))$ is 1 while in $r(X_{D_{2}})$ it is 0. Indeed, if $r(X_{D_{1}})$ is a subdiagram of $r(X_{D_{2}})$, we would have $r(X_{D_{1}})=r(X_{D_{2}})$ since the degree of $r(X_{D_{1}})$ is at least as big as that of $r(X_{D_{2}})$. This, however, is impossible.
@@
4. The diagram invariants are of finite type
@@
Choose a singular point $x$ of a singular doodle $D$ and let $v$ be the tangent vector to the last branch of $D$ at $x$. Write $v_{\perp}$ for the vector obtained from $v$ by rotating it by $\pi/2$ in the positive direction. Let $D^{\prime}$ and $D^{\prime\prime}$ be the singular doodles obtained by moving the last branch of $D$ slightly off $x$ in the directions $v_{\perp}$ and $-v_{\perp}$, respectively. A one-step resolution of $D$ at $x$ is the formal difference $D^{\prime}-D^{\prime\prime}$. This procedure can be iterated. The order of $D^{\prime}$ and $D^{\prime\prime}$ is one less than that of $D$; therefore, by taking consecutive one-step resolutions we arrive to a linear combination of usual doodles that we call the complete resolution of $D$.
@@
Since the complexity of each singular point is at least 2, the number $s$ of singular points is at most half the complexity of the singular doodle. If a singular doodle has complexity $c$ with $c>2n$, then $c>n+s$ and Proposition 4 implies that that $I_{n}$ is zero on it, which establishes Theorem 1.
Sketch of the proof of Proposition 4.
@@
A flat tangle $T$ gives rise to an arrow diagram whose skeleton is the 1-manifold parametrizing $T$. One may also talk about the sum of all the subdiagrams of the arrow diagram of $T$, we denote this sum by $I(T)$. In the same way as the singular doodle, we may define singular flat tangles. Proposition 4 is a consequence of the following fact: for the singular flat tangle $T_{k}$ with one singular point where $k$ branches of $T_{k}$ intersect, $I(T_{k})$ involves only diagrams with at least $k-1$ chords.
@@
5. Some remarks
1. In the definition of the spaces $\mathcal{A}_{n}$ and $\mathcal{Q}_{n}$, the coefficients of the diagrams can be taken in the field $\mathbb{F}_{2}$ with two elements. The diagram invariants defined in this fashion still distinguish doodles.
2. One may apply the constructions of this note to long doodles; these are doodles in $S^{2}$ that pass through a given point (thought of as the infinity in the Riemann sphere). Long doodles are parametrized by the real line and the preimages of the self-intersection points are naturally ordered. As a consequence, the branches of such a doodle at each self-intersection point are also ordered and each self-intersection point carries a sign which is positive whenever the ordered pair of tangent vectors to the doodle there forms a positive basis. Instead of arrow diagrams, long doodles may be encoded by chord diagrams with signed chords; the construction of the subdiagram invariant for such signed chord diagrams is entirely similar to the one described in the present note.

### 22 Feb 2024 v2

@@
Abstract
We define a complete invariant for doodles on a 2-sphere which takes values in a series of chord diagrams of a certain type. The coefficients in the diagrams with $n$ chords are finite type invariants of doodles of order at most $2n$.
keywords:
Doodle, finite type invariant, diagram invariant
[…]
To the memory of Sergei Duzhin
1 Introduction
The theory of finite-type knot invariants may be thought of as a nilpotency theory for knots. Nilpotent and residually nilpotent groups can be studied by embedding them into algebras of power series; most notably, the free group on $n$ letters $x_{1},\ldots,x_{n}$ maps into the algebra of the noncommutative power series in $X_{1},\ldots,X_{n}$ by
@@ 1 Introduction
The Kontsevich integral contains complete information about the finite type invariants of a knot. One may be tempted to construct universal finite-type knot invariants by using the same elementary method that works for the free groups. Namely, one can encode knots by diagrams of some kind and then assign to a knot a formal sum of the subdiagrams of its diagram. In particular, if knots are represented by so-called clasp diagrams [12], the formal sum of the subdiagrams taken by modulo diagrams with $n+1$ clasps is, in fact, a universal invariant of order $n$. This construction, however, is no replacement for the Kontsevich integral since the values of this invariant lie in a certain algebra which is rather more involved than the (already uncomputable) algebra of chord diagrams. Essentially, it is a restatement of the original approach of Vassiliev in different terms.
@@
In the present note, we consider the set of equivalence classes of curves in $S^{2}$ without triple self-intersections, called “doodles” by Khovanov [^1] The term “doodles” was first used [5] for a different class of curves. Khovanov’s terminology has now superseded that of Fenn and Taylor. in [9]. We show that the sums of subdiagrams give a finite type invariant of curves of doodles in $S^{2}$, which is complete in the sense that it distinguishes non-equivalent curves.
@@
2 Doodles and their arrow diagrams
@@
A doodle in $S^{2}$ is determined up to isotopy by its arrow diagram [^3] It would be more natural to use the term “Gauss diagram”. However, Arnold in [1] uses the term “Gauss diagram” for the underlying chord diagram of the arrow diagram, so we follow the terminology of Merkov. The term “arrow diagram” has also been used in a somewhat different context in [15]., which is a chord diagram with directed chords. The skeleton (that is, the outer circle) of the chord diagram corresponds to the parameter along the curve, the chords connect the inverse images of the crossings. At each crossing the branches of the doodle can be ordered in such a way that the tangent vectors to the first and the second branches form a positive basis of the tangent space to $S^{2}$. Each chord of the arrow diagram of the doodle is directed from the first branch to the second branch.
Figure 1: A doodle and its arrow diagram
@@
Note that by applying a move $R^{\prime}_{1}$ or $R^{\prime}_{2}$ in a way that decreases the number of crossings to a realizable arrow diagram we again obtain a realizable diagram. It follows that the minimal diagram equivalent to a realizable arrow diagram is also realizable. We arrive at the following
Proposition 2.1.
@@
A function on the set of isotopy classes of doodles that are preserved by the moves $R_{1},R_{2}$ is called an invariant of doodles in $S^{2}$. Proposition 2.1 implies that any function on the arrow diagrams that is preserved by the moves $R^{\prime}_{1},R^{\prime}_{2}$ gives rise to an invariant of doodles and that each doodle invariant arises in this way.
Write $\mathbf{D}$ for the set of all arrow diagrams and $\mathbf{Q}\mathbf{D}$ for the vector space over $\mathbf{Q}$ generated by this set. The moves $R^{\prime}_{1},R^{\prime}_{2}$ define a set of relations in $\mathbf{Q}\mathbf{D}$ which equate the left-hand side of a move to the right-hand side:
@@
the inverse limit of the $I_{i}$. Theorem 2.2 implies
Corollary 2.3.
@@
Let $V_{n}$ be the kernel of the projection $\mathcal{A}_{n+1}\to\mathcal{A}_{n}$. The space $\overline{\mathcal{A}}$ is isomorphic to the product of all the $V_{n}$. Another consequence of Theorem 2.2 is
Corollary 2.4.
@@
If a basis is chosen in each $V_{n}$, the map $\overline{I}$ becomes a series of arrow diagrams whose coefficients are finite type invariants, and the coefficient at any basis diagram in $V_{n}$ is a finite type invariant of order at most $2n$.
In fact, we will see that Theorem 2.2 holds even if $\mathcal{A}_{n}$ is replaced by a certain quotient of $\mathcal{A}_{n}$ which admits an explicit description. We consider it in the next section. In Section 4 we will show that $I_{n}$ is an invariant of order $2n$.
3 Quiver diagrams and the diagram invariants
@@
Figure 2: Reduction of a chord diagram
Lemma 3.1.
@@
Proof 3.2.
If a chord diagram $C$ has two pairs of adjacent chords, these pairs are either disjoint or have one chord in common. In both cases, the result of reducing $C$ with respect to both pairs does not depend on the order. It is obvious in the case when the pairs are disjoint; if they are not, inspection shows that the order of reductions does not matter in this case either.
Quiver diagrams [^4] The term “quiver” here has nothing to do with directed graphs and refers to a pack of arrows. are a generalization of arrow diagrams. A quiver diagram is a chord diagram in which the endpoints of the chords are labelled by nonnegative integers in such a way that the sum of the labels for every chord is at least 1. Quiver diagrams in which every chord has one end labelled with 1 and the other end with 0 may be identified with arrow diagrams if we take the direction of each chord from 0 towards 1. Define the degree of a quiver diagram to be the sum of its labels. The underlying chord diagram of a quiver diagram is obtained by forgetting the labels of the endpoints.
One may define the reduction of a quiver diagram with two adjacent chords as the reduction of the underlying chord diagram with the following rule for the labels: the labels of the adjacent endpoints of the adjacent chords are added together while the labels of the other chords remain the same, see Figure 3. We say that a quiver diagram is reduced if its underlying chord diagram is reduced. As in Lemma 3.1, any sequence of reductions of a given quiver diagram always terminates in the same reduced diagram.
Figure 3: Reduction of a quiver diagram
Let $\mathcal{Q}$ be the vector space over $\mathbf{Q}$ spanned by all the quiver diagrams modulo the subspace spanned by the following relations:
@@
Define $\mathcal{Q}_{n}$ to be the quotient of $\mathcal{Q}$ by the subspace spanned by the diagrams of degrees greater than $n$. For a reduced chord diagram $C$ without isolated chords and at most $n$ chords, let $\mathcal{Q}_{n}^{C}\subset\mathcal{Q}_{n}$ be the subspace spanned by the quiver diagrams whose underlying chord diagram is $C$.
Proposition 3.3.
@@
Proposition 3.4.
Given a reduced chord diagram $C$ without isolated chords, mark one endpoint for each chord. The set of all quiver diagrams of degree at most $n$ with the underlying chord diagram $C$ and the labels of the marked endpoints all zero form a basis of $\mathcal{Q}_{n}^{C}$.
@@
from where do we get
$D(p,q)=-D(p+1,q-1)+D(p+2,q-1)-D(p+3,q-1)+\ldots.$ (3)
@@
Rewriting in this way the relations in $\mathcal{Q}_{n}^{C}$ we obtain a tautology so all the diagrams whose marked endpoints are labelled with 0 are linearly independent.
@@
The second part of Theorem 2.2 is a consequence of the following statement:
Theorem 3.6.
@@
Figure 4: A pair of intersecting adjacent chords pointing in the same direction and what they might look like in a doodle. The points $B$ and $C$ must be joined in $S^{2}$ by a segment of the doodle containing no other intersection points, which is impossible.
Proof 3.7.
Assume that the arrow diagram $X_{D}$ of a minimal doodle $D$ contains a pair of adjacent chords. These chords must point in the same direction since otherwise they could be eliminated by a move $R_{2}^{\prime}$; this would mean that the doodle is not minimal. These chords cannot intersect each other, see Figure 4; the only possible situation when a pair of adjacent chords appears in a minimal doodle diagram is when these chords are parallel as shown in Figure 5 (a); in fact, in this case there may be a cluster with more than two parallel chords with adjacent endpoints as in Figure 5 (b). Such a cluster of parallel chords with adjacent endpoints divides the arrow diagram $X_{D}$ into two parts and no other chord intersects any of the chords of the cluster. The reduction of $X_{D}$ consists of replacing each maximal cluster of $m$ parallel arrows in $X_{D}$ with one chord whose initial point is marked by 00 and the final point by $m$. Denote the resulting reduced quiver diagram by $r(X_{D})$, see Figure 6. If $X_{D}$ is reduced to begin with, we set $r(X_{D})=X_{D}$.
Figure 5: Doodles whose arrow diagrams have adjacent chords
Figure 6: The diagrams $X_{D}$ and $r(X_{D})$.
Assume $D$ is minimal and has at most $n$ crossings. Since each chord in $r(X_{D})$ has one endpoint marked 0, we can choose a basis in $\mathcal{Q}_{n}$ as in Proposition 3.4 such that $r(X_{D})$ is a basis element. Then, the coefficient of $r(X_{D})$ in $q_{n}(I_{n}(D))$ is equal to 1 since there is precisely one subdiagram that gives $r(X_{D})$; namely, $X_{D}$ itself. It follows that $q_{n}(I_{n}(D))\neq 0$.
Now, let $D_{1}\neq D_{2}$ be two minimal doodles with at most $n$ crossings. Without loss of generality, assume that $D_{1}$ has at least as many crossings as $D_{2}$. Then, if the underlying chord diagrams of the arrow diagram $X_{D_{1}}$ and $X_{D_{2}}$ are different, $r(X_{D_{1}})\neq r(X_{D_{2}})$ and we can choose a basis for $\mathcal{Q}_{n}$ in such a way that both $r(X_{D_{1}})$ and $r(X_{D_{2}})$ are basis elements. Then, the coefficient of $r(X_{D_{1}})$ in $q_{n}(I_{n}(D_{1}))$ is 1 while in $r(X_{D_{2}})$ it is 0. Indeed, if $r(X_{D_{1}})$ is a subdiagram of $r(X_{D_{2}})$, we would have $r(X_{D_{1}})=r(X_{D_{2}})$ since the degree of $r(X_{D_{1}})$ is at least as big as that of $r(X_{D_{2}})$. This, however, is impossible.
@@
4 The diagram invariants are of finite type
@@
Choose a singular point $x$ of a singular doodle $D$ and let $v$ be the tangent vector to the last branch of $D$ at $x$. Write $v_{\perp}$ for the vector obtained from $v$ by rotating it by $\pi/2$ in the positive direction. Let $D^{\prime}$ and $D^{\prime\prime}$ be the singular doodles obtained by moving the last branch of $D$ slightly off $x$ in the directions $v_{\perp}$ and $-v_{\perp}$, respectively. A one-step resolution of $D$ at $x$ is the formal difference $D^{\prime}-D^{\prime\prime}$. This procedure can be iterated. The order of $D^{\prime}$ and $D^{\prime\prime}$ is one less than that of $D$; therefore, by taking consecutive one-step resolutions we arrive at a linear combination of usual doodles that we call the complete resolution of $D$.
@@
Since the complexity of each singular point is at least 2, the number $s$ of singular points is at most half the complexity of the singular doodle. If a singular doodle has complexity $c$ with $c>2n$, then $c>n+s$ and Proposition 4.1 implies that that $I_{n}$ is zero on it, which establishes Theorem 2.2.
Proof 4.2 (Sketch of the proof of Proposition 4.1).
@@
A flat tangle $T$ gives rise to an arrow diagram whose skeleton is the 1-manifold paramet-rizing $T$. One may also talk about the sum of all the subdiagrams of the arrow diagram of $T$, we denote this sum by $I(T)$. In the same way as the singular doodle, we may define singular flat tangles. Proposition 4.1 is a consequence of the following fact: for the singular flat tangle $T_{k}$ with one singular point where $k$ branches of $T_{k}$ intersect, $I(T_{k})$ involves only diagrams with at least $k-1$ chords.
@@
5 Some remarks
5.1 In the definition of the spaces $\mathcal{A}_{n}$ and $\mathcal{Q}_{n}$, the coefficients of the diagrams can be taken in the field $\mathbb{F}_{2}$ with two elements. The diagram invariants defined in this fashion still distinguish doodles.
5.2 One may apply the constructions of this note to long doodles; these are doodles in $S^{2}$ that pass through a given point (thought of as the infinity in the Riemann sphere). Long doodles are parametrized by the real line and the preimages of the self-intersection points are naturally ordered. As a consequence, the branches of such a doodle at each self-intersection point are also ordered and each self-intersection point carries a sign which is positive whenever the ordered pair of tangent vectors to the doodle forms a positive basis. Instead of arrow diagrams, long doodles may be encoded by chord diagrams with signed chords; the construction of the subdiagram invariant for such signed chord diagrams is entirely similar to the one described in the present note.